OpenStudy (anonymous):

What is the difference of (5x - 2/ 4x) - (x - 2/ 4x)?

OpenStudy (imtant):

So it will be like this 5x-2-(x-2)/(4x) 4x/4x x=1

OpenStudy (imtant):

the denominator is same 4x so you just need to write once.

OpenStudy (anonymous):

How did you do that? Show me every step please.

OpenStudy (anonymous):

Simplify (5x-2/4x) to 4 2/4x (or 4.5x), then simplify (x-2/4x) to 2/4x (or .5x). Next you subtract 4.5x- .5x= 4x. Substitute 1 for x to get your answer.

OpenStudy (imtant):

5x-2-(x-2)=0 5x-2-x+2=0 So then 5x-x-2+2=0 It will be 4x.. while -2+2=0 Lastly 4x/4x=1 For example:2/2=1,right?ok?

OpenStudy (anonymous):

What is the sum of 1/2b + b/2?

OpenStudy (anonymous):

@imtant

OpenStudy (imtant):

Ok this ques. you need to multiply b/2 with b 1/2b + b^2/2b so it is b^2+1 all over by 2b.

OpenStudy (anonymous):

What's the difference of (r + 2/ r + 4) - (3 / r + 1)?

OpenStudy (anonymous):

@imtant

OpenStudy (anonymous):

I need help with this question and just one more.

OpenStudy (imtant):

Ok,wait.

OpenStudy (imtant):

Multiply crossly.(r+2)(r+1)-(3)(r+4) divide all over by (r+4)(r+1) it will be r^2+r+2r+2-(3r+12) divide all by (r+4)(r+1) so,r^2+3r+2-3r-12 all divide by (r+4)(r+1) then r^2-10 divide all by (r+4)(r+1)

OpenStudy (anonymous):

Kathal in drove 180 miles on her vacation. She drove an average of 1.5 times faster on the second 90 miles of her trip than she did on the first 90 miles of her trip. Which expression represents the time she spent driving? Let x = her speed on the first half of the trip.

OpenStudy (anonymous):

*Katalin*

OpenStudy (anonymous):

@imtant

OpenStudy (anonymous):

Hellooooo???

OpenStudy (anonymous):

:(

OpenStudy (imtant):

1 trip=90 x miles 2 trips= 90x times 1.5=135 x miles So the time is 180=135 x x=1.33 for time.

OpenStudy (anonymous):

These are the answer choices: A. 150/x B. 300/x C. 150x D. 225/x

OpenStudy (anonymous):

(5x-2/4x)-(x-2/4x) =(20x^2-2)/4x-(4x^2-2)/4x =(20x^2-2-4x^2+2)/4x =16x^2 /4x =4x

OpenStudy (anonymous):

? @imtant @surjithayer

OpenStudy (anonymous):

Katalin drove 180 miles on her vacation. She drove an average of 1.5 times faster on the second 90 miles of her trip than she did on the first 90 miles of her trip. Which expression represents the time she spent driving? Let x = her speed on the first half of the trip. A. 150/x B. 300/x C. 150x D. 225/x